Single Plantwide Factory Overhead Rate Salty Sensations Snacks Company manufactures three types of snack foods: tortilla chips, potato chips, and pretzels. The company has budgeted the following costs for the upcoming period: Factory depreciation $14,315 Indirect labor 35,476 Factory electricity 4,045 Indirect materials 8,402 Selling expenses 19,916 Administrative expenses 11,203 Total costs $93,357 Factory overhead is allocated to the three products on the basis of processing hours. The products had the following production budget and processing hours per case: Budgeted Volume (Cases) Processing Hours Per Case Tortilla chips 3,900 0.12 Potato chips 5,700 0.10 Pretzels 2,100 0.15 Total 11,700 If required, round all per-case answers to the nearest cent. a. Determine the single plantwide factory overhead rate. $ __ per processing hour b. Use the factory overhead rate in (a) to determine the amount of total and per-case factory overhead allocated to each of the three products under generally accepted accounting principles.
Answer #1

A B C D w N 5 1 a) Processing hours per Total Processing Budgeted volume 2 case hours 3 Tortilla chips 3,900 0.12 468 4 Potat

D Total Processing hours =+B3*C3 =+B4*C4 =+B5*C5 =+D3+D4+D5 5 А B 1 a) Processing hours per Budgeted volume 2 case 3 Tortilla